OpenStudy (1dfanatic):

-2/3+4/6 (fractions simplify)

OpenStudy (muscrat123):

common denominator

OpenStudy (muscrat123):

-2/3 = -4/6 b/c -2/3 times 2 equals -4/6

OpenStudy (1dfanatic):

so do i multiply both the fractions by 2? @muscrat123

OpenStudy (muscrat123):

only one

OpenStudy (muscrat123):

b/c u want common denominators

OpenStudy (shreenithaa):

-2/3 +4/6 u have to have the same denominator in order to be able to add both (-2/3)=-4/6 -4/6+4/6 what the answer?

OpenStudy (1dfanatic):

0?

OpenStudy (anonymous):

it is zero cuz -2/3 is the same as -4/6 so -2/3 +4/6 is the same as -2/3 +2/3 and vice versa

OpenStudy (shreenithaa):

think so... yeah

OpenStudy (anonymous):

Yup, your final answer is 0

OpenStudy (1dfanatic):

and the bottom is 18 right?

OpenStudy (1dfanatic):

because it's a fraction

OpenStudy (shreenithaa):

nonono -2/3 equals -4/6 right u multiply both the denominator and the numerator by two only to -2/3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is 0/18

OpenStudy (shreenithaa):

how?

OpenStudy (anonymous):

me??? @shreenithaa

OpenStudy (shreenithaa):

yeah... lol

OpenStudy (muscrat123):

0/18 is equal to 0

OpenStudy (shreenithaa):

yup @muscrat123 is right

OpenStudy (shreenithaa):

but i just done get how u guys got 18

OpenStudy (muscrat123):

me eitheer

OpenStudy (shreenithaa):

what grade r u in and state

OpenStudy (anonymous):

Its the smallest dcommon denominator that both the denominators have

OpenStudy (muscrat123):

i just know 0/18 is 0, so it doesnt matter what that second number is. and me? @shreenithaa

OpenStudy (shreenithaa):

nope its 6 3x2=6 6x1=6

OpenStudy (anonymous):

Im in 10th grade and taking Algebra 2 and I live in SC

OpenStudy (shreenithaa):

8th california

OpenStudy (shreenithaa):

algebra 1

OpenStudy (muscrat123):

im in 7th grade, with 9th grade classes and im taking Algebra 1 in Florida

OpenStudy (1dfanatic):

9th algebra 1 and pa

OpenStudy (anonymous):

tru tru @shreenithaa sorry i missed that thanks so the answer would be 0/6 let me check my online calculator

OpenStudy (shreenithaa):

its k

OpenStudy (anonymous):

This is what i always use and it says the answer is 0/18 http://www.calculatorsoup.com/calculators/math/fractions.php

OpenStudy (shreenithaa):

maybe it a type of calculator that multiplies the denominators

OpenStudy (shreenithaa):

anyways everyones right lol

OpenStudy (shreenithaa):

bye